OpenStudy (anonymous):

what is 17/20 0.92 80% 6/8 in least to greatest

OpenStudy (anonymous):

6/8,80%,17/20,0.92

OpenStudy (anonymous):

17/20=0.85 80%=80/100=0.80 6/8=3/4=0.75 so now you can arrange them in order